A greyish metal rod with no markings is found in the laboratory and needs to be identified. A student performs coffee-cup calorimetry to determine the specific heat of the metal. The metal (7.00g) is first heated to 90 degrees Celsius, and then put into 20.0 g of water at 43.6 degrees Celsius. The temperature after equilibration is 44.5 degrees Celsius. What is the identity of the metal?

metal specific heat (J.gC)
aluminum 0.900
silver 0.236
lead 0.128
iron 0.450
1
Expert's answer
2017-03-08T08:47:06-0500
Q(H₂O)=cm(t-t₁)=4200 J/kg*ᵒC*0.02kg*(44.5-43.6)=75.5J

Q(Me)=cm(t-t₁)=c*0.007kg*(44.5- 90)=-0.3185cJ

Q(H₂O)=Q(Me)

0.3185c=75.6

c=237.36

c(Me)=236.36 J/kg*ᵒC=0.236J/gᵒC

silver 0.236